OpenStudy (anonymous):

A geosyncchronous equatorial orbiting (GEO) satellite orbits 22.300 miles above the equator of Earth. It completes one full revolution each 24 hours. Assume Earth's radius is 3960 miles.

OpenStudy (anonymous):

Let me try......ok, I can assume that :)

OpenStudy (anonymous):

Done, I have assumed that the Earth's radius is 3960 miles. Now what?

OpenStudy (anonymous):

a) How far will the GEO satellite travel in one day ?

OpenStudy (anonymous):

b) Whay is the satellite's linear velocity in miles per hour ?

OpenStudy (anonymous):

Then it travels in a circle with radius 3960+22.3 miles in a day so 2πr with r=3988.3.

OpenStudy (anonymous):

For part a. about 25,046.524 miles

OpenStudy (anonymous):

For part b. just divide by 24 and you get 1043.60517 miles per hour.

OpenStudy (anonymous):

a) 1 day = 3960+ 22.982.3 than ?

OpenStudy (anonymous):

3960 + 22.982.3 represents the outer radius because satellite orbits above equator of the earth

OpenStudy (anonymous):

The to solve for how far it will travel in one day (24 hours) you use the circumference formula: 2*pi*r...r = 3988.3

OpenStudy (anonymous):

So 2*(pi)*(3988.3) = 25,046.524 miles

OpenStudy (anonymous):

Then linear velocity in miles per hour is 25,046.525 miles/ 24 hours = 1043.6052 miles / hour

OpenStudy (anonymous):



OpenStudy (anonymous):

radius small circle = 3960 big circle= 33.300?

OpenStudy (anonymous):

no radius of big circle is 3960 + 33.300 = 3993

OpenStudy (anonymous):

Think about it you start from center of inside circle

OpenStudy (anonymous):

than small is ?

OpenStudy (anonymous):

small is what you were given to start with: 3960

OpenStudy (anonymous):

full revolution each 24 hours, for travel big circle ?

OpenStudy (anonymous):

Good full revolution of 24 hours is 2*pi*r pi is constant = 3.14..or use your calculator for approximation r = 3993

OpenStudy (anonymous):

Do you get it?

OpenStudy (anonymous):

22.300+3960=26.260

OpenStudy (anonymous):

26.260^2*3.14?

OpenStudy (anonymous):

I don't know how u get 3993?